OpenStudy (anonymous):

Ellen wants to purchase a book that is regularly priced at $18. The book is discounted 15%. She also needs to pay a 6% sales tax on the discounted price. What is the total amount Ellen will pay for the book? Enter your answer in the box.

OpenStudy (anonymous):

please dont give answer. Just help me figure it out

OpenStudy (texaschic101):

regular price = 18 discount price is 15% of 18.....so for the discount , we turn percent to decimal, then multiply discount = 18 * 0.15 = ??

OpenStudy (anonymous):

2.7

OpenStudy (texaschic101):

correct...now thats the discount....for the discount price, we take original price minus discount...... discount price = 18 - 2.70 = ??

OpenStudy (anonymous):

15.3?

OpenStudy (texaschic101):

correct...thats the discount price now we need to add 6% sales tax first we find the sales tax by taking 6% of 15.30......turn percent to decimal, then multiply. 15.30 * .06 = ? (this is your sales tax)

OpenStudy (anonymous):

.918 also why isnt sales tax .6? why the 0?

OpenStudy (texaschic101):

0.918 is correct.....you can probably round it to 0.92. 6% tax....we turn it into a decimal by dividing by 100 = .06

OpenStudy (texaschic101):

so we have discount price = 15.30 and sales tax = 0.92 for the total price, we add them... 15.30 + 0.92 = ? (total price)

OpenStudy (anonymous):

16.22..... thats the answer?

OpenStudy (texaschic101):

that is the answer :)

OpenStudy (anonymous):

THANK YOU SO MUCH!

OpenStudy (texaschic101):

anytime :)